A coin is thrown 3 times. List out the sample space. What is the probability that at least one head is obtained?


1
Expert's answer
2021-11-08T16:12:05-0500

When you flip a coin 3 times, then there are "2^3=8" possible outcomes.


"S=\{HHH, HHT, HTH, THH,"

"HTT, THT, TTH, TTT\}"

"P(at\ least\ 1\ H)=1-P(3\ T)=1-\dfrac{1}{8}=\dfrac{7}{8}"